### v3

# Case Law Grounding: Using Precedents to Align Decision-Making for Humans and AI

Abstract.
Communities and groups often need to make decisions grounded by social norms and preferences, such as when moderating content or providing judgments for aligning AI systems. Prevailing approaches to provide this grounding have primarily centered around constructing high-level guidelines and criteria, similar to legal “constitutions”. However, it can be challenging to specify social norms and preferences consistently and accurately through constitutions alone. In this work, we take inspiration from legal systems and introduce “case law grounding” (CLG)—a novel approach for grounding decision-making that uses past cases and decisions (precedents) to ground future decisions in a way that can be utilized by human-led processes or implemented through prompting large language models (LLMs). We evaluate how accurately CLG grounds decisions with five groups and communities spread across two decision task domains, comparing against a traditional constitutional grounding approach, and find that in 4 out of 5 groups, decisions produced with CLG were significantly more accurately aligned to ground truth: 16.0–23.3 %-points higher accuracy using the human-led process, and 20.8–32.9 %-points higher when prompting LLMs. We also evaluate the impact of different configurations of CLG, such as the case retrieval window size and whether to enforce binding decisions based on selected precedents, showing support for using binding decisions and preferring larger retrieval windows. Finally, we discuss the limitations of our case-based approach as well as how it may be best used to augment existing constitutional approaches when it comes to aligning human and AI decisions.
alignment; human judgment; large language models; subjectivity; consistency; content moderation; policy; decision-making
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Many technological systems today are entangled in social decision-making situations. Online spaces for inter-personal communication, like social media platforms, need to contend with decisions around how to moderate content in a way that maintains a healthy community (Gongane et al., 2022). General purpose AI assistants, like LLM-backed chatbots, also need to contend with decisions around how to generate answers in a way that users find appropriate (Ji et al., 2023; Shen et al., 2023; Gabriel, 2020). The scale of these systems means that getting the decisions wrong can have major negative consequences for people (Haimson et al., 2021; Weidinger et al., 2021).
Unlike systems that work with more objective determinations, decisions around social concepts are especially hard because: (1) between groups or communities, there often isn’t a shared consensus around what is “correct”, and (2) within a group or community, norms and expectations are often implied and not formalized. This makes it hard to make consistently good decisions at scale—groups and communities either make efforts to ensure decision quality through highly involved procedures (e.g., deliberation (Im et al., 2018; Small et al., 2021), digital juries (Fan and Zhang, 2020), expert panels (Klonick, 2020)), or they might trade off some decision quality for scale (e.g., building classifiers and auto-moderation (Jhaver et al., 2019b; Wulczyn et al., 2017), outsourcing to paid crowd workers (Satariano and Isaac, 2021)).
One way that current systems have attempted to bridge this demand has been to try and distill social norms and preferences into general decision criteria. For example, moderators of communities on social media platforms often create community guidelines that establish what content is allowable, and how violations are to be dealt with (Chandrasekharan et al., 2018a; Chandrasekharan and Gilbert, 2019). Similarly, builders and operators of AI assistant services often come up with principles (sometimes with help from the general public (Huang et al., 2024)), that are then used to synthesize training and fine-tuning examples (Bai et al., 2022) or become part of the prompts (Lin et al., 2024) for models. However, in practice, imprecise language can leave room for ambiguity (Fiesler et al., 2018; Chen et al., 2021) around edge cases, and varied decision-makers can result in divergent interpretations (Gordon et al., 2021). Is there another way we could better support making socially-aligned decisions at scale? For a solution, we look to the realm of legal systems: In legal systems, the creation of principles and guidelines to ground decisions parallels that of creating and applying constitutional law. However, many legal systems also utilize a different way to ground more nuanced decisions—case law. With case law, rather than relying on lawmakers (or, in our case, moderators and experts) to resolve every edge case and ambiguity through modifications and additions to the constitution, decision-makers can also rely on past decisions as sources of clarification and disambiguation.
Taking inspiration from case law legal systems, we introduce the case law grounding (CLG), a general workflow for grounding decisions on subjective or socially-aligned tasks that can be support human-led or LLM prompting processes. Like with case law in the legal realm, CLG focuses on making use of precedents (past decisions) as the primary source of guidance for new decisions. CLG accomplishes this through 3 steps: (1) case retrieval, where past cases potentially relevant to the current decision are retrieved; (2) precedent selection, where retrieved precedents are further refined for applicability to the current decision; and finally (3) decision synthesis, where the selected precedents and decisions are aggregated and synthesized into a decision for the new case. As online groups and communities often need to make these types of decisions at-scale, we also examine both ways in which the CLG workflow could be implemented: the human-led version targeting use cases where members within the group or community are the decision-makers; and an LLM-prompting based version targeting use cases where groups and communities may want to scale up with the help of AI while also ensuring decisions remain aligned to their own values.
Recruiting participants from 3 Reddit communities and 2 demographically-recruited population groups, we evaluated both human against a traditional workflow that uses constitutional grounding to support decisions. In the human-led process, participants made binary “keep”/“remove” decisions around content moderation outcomes (Park et al., 2021) within a community they frequent, or rated toxicity of social media posts (Kumar et al., 2021) for their demographic group, recruiting $n=35$ Redditors and $n=60$ crowd workers for the mod and tox domains respectively. We also evaluated an automated version of CLG for the same task domains, implemented through prompting an LLM.
In summary, in this work we make the following contributions:
  • We introduce a general framework for supporting group-aligned decision-making in subjective and social contexts, Case Law Grounding (CLG), and detail how CLG can be applied to two decision-making situations: a human-led process where community and group members make decisions grounded by precedents, and an automated process where an LLM is prompted to evaluate precedents instead of human annotators.
  • We evaluate CLG with participants drawn from 5 groups (based on communities or demographic population groups) across two decision task domains, comparing against a more traditional approach to grounding decisions via constitutional rules and guidelines. We find that:
    • RQ1: When using either the human-led or LLM prompting process, CLG was able to produce final decisions that more accurately matched observed ground truth compared to decisions grounded with constitutions, with statistically significant improvements observed in 4 out of 5 groups.
    • RQ2a: Examining the retrieval process by simulating smaller retrieval window sizes, we find that for most domains, larger retrieval windows resulted in higher accuracy, however scaling characteristics varied across groups.
    • RQ2b: Comparing setups where precedents were binding versus those where they only served as examples, we find that the binding nature of precedents is important—in the 4 groups from RQ1 where the binding configuration produced statistically significant improvements in decision accuracy, the non-binding version did not produce statistically significant differences.
    • RQ3: We find that beyond the human-led process, LLMs were able to scale up to larger datasets, while still producing statistically significant improvements across all 6 tested groups.
  • Finally, we discuss limitations of human-led decision-making with case law grounding that arise from human evaluation of concrete cases and discuss how constitutional and case law grounding can complement each other in supporting group-aligned subjective decision-making.
Figure 1. A diagram contrasting different approaches to align decisions: (1) traditional “constitutional” grounding design, involving applying abstract guidelines to construct judgments; and (2) “case law” grounding proposed in this work, that utilizes retrieval of past cases followed by selection of applicable precedents to construct judgments.
[…]
A diagram illustrating two types of ways to ground decisions. On the top is a constitutional workflow, which follows a process of referring to guidelines, applying high-level rules, and then making a final judgment. Below is a case law workflow, which follows a process of retrieving relevant cases, selecting locally-relevant precedents, and then automatically deriving a final judgment. The second step in each process is highlighted purple indicating it forms the core of what is an alignment mechanism.
2. Related Work
In this section, we discuss how our work relates to prior work around decision-making in social or subjective settings, constitutional approaches for consistent decision-making, and precedents of using cases to aid decision-making in prior social and technical systems.
2.1. Constitutions for Social and Subjective Concepts
The rise of online platforms has led to an increased demand for making socially-situated decisions through technological systems. For example, online communities such as Wikipedia, Reddit, and Facebook all need to make moderation decisions around what content is objectionable and should be removed (Van Aken et al., 2018; Kumar et al., 2021; Bhuiyan et al., 2020; Ross et al., 2017). However, the definitions of these concepts, like the criteria for moderation, can differ significantly between groups and communities (Chandrasekharan and Gilbert, 2019; Chandrasekharan et al., 2018a), which means it is important to ground decisions relevant to these groups locally to each group’s own preferences rather than universally across the entire platform.
2.1.1. Challenges of Creating Constitutions
Constitutions as Guidelines: Smaller online communities, such as subreddits, forums, or Discord servers, often rely on volunteer moderators drawn from members of the community to manually make moderation decisions (Matias, 2019). In these cases, moderators might already start with an intuitive understanding of what the criteria are simply based on their experience in the community. However, the socially-constructed intuition of what is desirable might not be aligned across all community members (Weld et al., 2021), and can lead to user frustration around decisions (Jhaver et al., 2019a). Community guidelines can codify these intuitive preferences as a constitution, which can then communicate criteria around moderation to all members.
However, in practice, moderators often have additional criteria undocumented in community guidelines, leading to mismatches when compared against real decisions (Juneja et al., 2020). It can also be difficult to neatly translate some concepts into actionable criteria than others, leading to guidelines that might still leave some concepts ambiguous and open to interpretation (Arhin et al., 2021; Fiesler et al., 2018). Finally, the high amount of effort that must be put into making clear and self-consistent guidelines (Im et al., 2018) leads to practical limits for many communities.
Constitutions as Rubrics: On bigger platforms that host content for millions or billions of people, consistency and transparency plays a more paramount role (Klonick, 2020). In these cases, organizations can easily justify effort to develop complex constitutions and even costs around training paid moderators to execute decisions based on the constitution (Roberts, 2014; Satariano and Isaac, 2021). One leaked set of internal manuals from Facebook contained over 1,400 pages across over 100 manuals (Guardian, 2017).
However, the extensive iteration often means that rubric-style constitutional grounding, while efficient, can be inflexible to context (Guardian, 2017; Press, 2024), which is especially undesirable given the social nature of such decisions. In cases where there are cultural or even sub-community level differences, it may be intractable to accommodate such nuances with rubric-style constitutional grounding (Jiang et al., 2021). When constitutions are relied on as the main way to guarantee consistent decisions, the resulting constitutions often become laundry lists of positive and negative carve-outs and examples for each category of content (Klonick, 2017).
Constitutions as Code and Models: Code and machine learning models can also be integrated into constitutional grounding when they are built with the intent of capturing some abstract aspect of the criteria. For example, classifiers can help identify and tag concepts (Wulczyn et al., 2017; Müller-Birn et al., 2013), which can be combined with constitutional guidelines to make automated decisions. Additionally, communities can build code-based constitutions to automatically execute decisions through systems like AutoModerator (Jhaver et al., 2019b), ORES (Halfaker and Geiger, 2019), and PolicyKit (Zhang et al., 2020).
However, building constitutions as code and models depends on accurate measurements of concepts that are often contextualized only in social settings and lack objective definitions. These types of annotations by nature come with annotator positionality (Chandrasekharan and Gilbert, 2019; Chandrasekharan et al., 2018a; Arhin et al., 2021; Denton et al., 2021) that result in conflicting or unclear decisions. The data being labeled (Chen et al., 2021) or annotation guidelines (Chang et al., 2017) can also be incomplete or ambiguous. Attempts to create objective instructions for annotation have also seen researchers and domain experts disagreeing on definitions (Fortuna et al., 2020) or criteria (Bhuiyan et al., 2020).
More recently, the advent of prompting-based control of large language models (LLMs) has also given rise to natural language constitutions being directly executed through system prompts in a zero-shot or few-shot way (Lin et al., 2024; Han, 2023). However, the limited control over how these models apply prompts means it can be difficult to build robust prompts that appropriately balance competing community desiderata (Raghavan, 2024).
2.1.2. Challenges of Using Constitutions for Decision-Making
Constitutions (in the form of rubrics and guidelines) have also been used extensively by crowdsourcing systems to improve annotation quality (Wu and Quinn, 2017), and thus many of the major improvements on making constitutional grounding more consistent can be found in systems for instructing the crowd.
One approach for improving consistency has been to continuously identify and refine problematic parts of rubrics associated with inconsistency, either automatically (Pradhan et al., 2022) or with worker cooperation (Bragg and Weld, 2018). Alternatively, other work has focused on real time training and evaluation during the decision-making phase of annotation, either in the form of gating (Liu et al., 2016) at the start, continuous evaluation during the task (Hettiachchi et al., 2021), or dynamic training (Werling et al., 2015).
However, applications of rubrics for consistency has been most beneficial in the area of more objective annotation tasks (Snow et al., 2008; Rashtchian et al., 2010). For primarily subjective surveys (Bentley et al., 2017) and annotations over socially-constructed concepts (Chen and Zhang, 2023), constructing comprehensive rubrics may be counterproductive (removing the subjectivity we would like to capture) or self-defeating (when we must come up with criteria that we are collecting judgments to define in the first place).
Because constitutional guidelines are created by humans, they can often omit aspects considered shared knowledge (Chandrasekharan et al., 2018b) in a community or seen as “common sense”, or fail to document criteria being applied in practice (Juneja et al., 2020). This can make it challenging for anyone without such context to accurately judge situations following guidelines (Jhaver et al., 2019a, c). Additionally, as norms and preferences shift over time (Lin et al., 2017), decisions based on past implicit knowledge may silently become out-of-date without easy means of correction. Together, this makes direct adaptation of constitutions to automated systems potentially fragile, leading to unexpected behaviors when the guidelines are under-defined or conflicting (Gordon et al., 2021).
Just like human decision-makers who may not always consistently and accurately interpret and decide based on constitutions, LLMs can fail in this respect as well, with prior work observing that LLMs grounded using constitutions can perform better than random for certain communities (Kumar et al., 2024).
2.2. Adopting Cases to Aid Judgments
The use of cases as judgment aids has been adopted by a wide variety of social and technological systems in the past. While sets of cases are not as compact or easily interpretable at a high level as constitutional policies, they can contribute to high-quality judgments by provide two types of support that complement deficiencies of constitutions: (1) grounding complex concepts—cases are specific and concrete and thus can illustrate concepts when abstract language fail; and (2) serving as a fine-grained policy definition—sets of cases with associated decisions can be adapted as a means to define fine-grained community- or group-specific policy via retrieval and interpolation.
2.2.1. Examples of Using Cases to Ground Complex Concepts
Cases and case studies have broadly been used as a medium for reasoning around complex scenarios in various fields similar to the socially-constructed judgments we focus on. For example, moral case deliberation (Pierce, 2013; Sush and Najdowski, 2021; Emelin et al., 2021) is often used in medical settings to reflect on ethical situations (Molewijk et al., 2008), and case studies are common in business educational settings too (Farashahi and Tajeddin, 2018). In legal education, fictional cases are frequently used in exercises like moot court (Gaubatz, 1981; Ringel, 2004), as a medium for engaging in deliberative reasoning. Case-based reasoning (CBR) that focuses on analyzing the circumstances around specific case scenarios also has a long and rich history as a method for thinking about moral philosophy (Paulo, 2015) and in the realm of jurisprudence in the US legal system (Kagan, 2019).
Examples of input items and resulting judgments, often used to clarify rubrics and guidelines, can also be seen as a way of utilizing cases for grounding. In crowdsourced annotation, examples can augment text in rubrics to demonstrate or contrast how judgments should be made (Wu and Quinn, 2017), act as as tests to evaluate whether a concept or criteria was accurately understood (Liu et al., 2016; Hettiachchi et al., 2021; Singla et al., 2014), and even pose as reference anchors to ground the levels on a subjective scale (Chen et al., 2021).
Finally, the detail provided by individual cases means that they can act as a medium for deliberation to reconcile differences and work towards shared understanding. We can see example of this in legal realm around adjudication (Carpenter, 1917), medical realm for diagnosis (Bertens et al., 2013; Schaekermann et al., 2019), in communities for working out moderation policy (Fan and Zhang, 2020; Pan et al., 2022), and more generally just around challenging tasks involving different lines of reasoning (Chen et al., 2019).
2.2.2. Case-based Reasoning and Data-centric Approaches to Defining Concepts
Case-based reasoning (CBR) has been applied in computational areas as well, forming a basis for computational reasoning systems in some classical AI systems (Aamodt and Plaza, 1994). Here, CBR is formalized as a 4-step process: (1) retrieve: where relevant “cases” (such as past problems and solutions) to a target problem are recalled; (2) reuse: where the reasoning and solutions from the retrieved cases are mapped to the new problem; (3) revise: where solutions are tested and any errors revised; and (4) retain: where the (revised) solution and problem are stored for use in future retrieval. However, application of CBR on more subjective tasks that require interpreting natural language has been limited due to limited capabilities of earlier models that focused on addressing objective reasoning tasks through formalized and symbolic approaches around what constitutes arguments and reasoning  (Ashley, 1989; Craw and Aamodt, 2018; Craw et al., 2001). In this work, we draw from general concepts in CBR around the process for making decisions with cases, while also focusing on adapting these ideas to subjective decision-making by (often non-expert) community members around socially-situated decisions based on norms and preferences which can be difficult (and possibly undesirable) to create symbolic formalizations of.
Viewed through a machine learning lens, cases also arise naturally as a data-centric view to define human criteria, including those that are subjective or dependent on social context (Simons et al., 2020; Jiang and de Marneffe, 2022). The training of classifiers using human annotated data is itself a form of grounding (a model’s) decisions through the medium of cases. However, viewing training data as an example of cases also shows the limitations of how we work with training data. For one, annotated datasets on socially-situated or subjective concepts are often produced by annotators who don’t have the correct context to judge, leading to biased results (Arhin et al., 2021; Denton et al., 2021). Additionally, aggregation used during pre-processing can hide disagreements that would have required human reconciliation (Gordon et al., 2021, 2022). Still, when data is relatively abundant, machine learning approaches tend to outperform expert constructed systems, though the best approaches depend on a combination of expert and learned aspects (Gennatas et al., 2020).
2.2.3. LLM Prompting as a Parallel of Human Judgment
Recent advances in natural language processing have brought us large language models (LLMs), which can act as general purpose problem solving systems that can be interacted with natural language prompts (Shi et al., 2023; Shen et al., 2023). The increasing adoption of LLMs in sociotechnical systems has presented both challenges and opportunities for decision-making at scale in social settings.
In recent works, LLMs have been adapted to complete a variety of complex reasoning and decision tasks (Minaee et al., 2024; Taori et al., 2023). Unlike traditional machine learning classification, where models produce decisions opaquely in a way that would differ significantly from a precedent reasoning process done by humans, the new capability of being able to work with natural language instructions and inputs makes it possible to create a close parallel of the human decision-making process through the use of LLM prompting. Along these lines, prompting strategies like Chain-of-Thought prompting (Wei et al., 2022) and prompt chains (Wu et al., 2021) have been proposed to improve the quality of model outputs that involve reasoning around a task. While other works have revisited ideas from crowdsourcing, breaking larger tasks into sub-tasks and applying quality control feedback to improve LLM capabilities (Wu et al., 2023; Grunde-McLaughlin et al., 2023). There is even opportunity for LLM-backed agents to reproduce group- or community-level social insights, with some works examining cultural patterns by interrogating LLM outputs, while others use LLMs to simulate samples of populations via prompting models with “personas” (Aher et al., 2022; Cheng et al., 2023). These advances coupled with demand from communities to scale up decision-making inspired us to additionally evaluate the CLG workflow when LLM prompts are constructed to take the place of human annotators in the human-led process.
However, while LLMs may be able to capture some aspects of social context, they are also susceptible to biases with simulated personas behaving in overtly stereotypical ways (Gupta et al., 2024). To address some of the deficiencies of LLMs, retrieval augmented generation (Lewis et al., 2020b, a) (RAG) has been proposed which combine retrieval from knowledge bases of facts with reasoning capabilities of LLMs and prompt chains, to reduce model reliance on its own knowledge distributions, which may be biased, inaccurate, or out-of-date. While most applications of RAG have been centered around (factual) knowledge retrieval to combat hallucinations (Béchard and Ayala, 2024; Lewis et al., 2020a), recent work has also looked at retrieval over social contexts like personalization (Baek et al., 2024). The increasing recognition of the value of dynamic retrieval of past examples to introduce external factual knowledge in LLM space also nicely parallels ideas in CLG where precedents are retrieved to introduce knowledge around group and community norms and values to ground subjective socially-situated decisions.
Finally, while prompting has also been applied to the task of aligning the models themselves, through approaches like in-context learning (Han, 2023; Hendrycks et al., 2020) which incorporate static few-shot examples as a part of system prompts to guide model behavior, how models apply this information towards generated output can be opaque and unreliable (Arditi et al., 2024). Applying precedents this way would potentially run into similar issues around consistency as constitutional approaches with human-led decision-making. Thus, rather than rely on the models to correctly synthesize decisions by itself through examples, we implement the LLM-prompting version of CLG in a way that follows the human-led version—limiting model decisions to selection of precedents rather than and synthesizing the final decision directly.
3. Case Law Grounding
In this work, we introduce and explore the framework of Case Law Grounding (CLG), a novel approach to supporting decision-making over subjective and socially-situated decisions aligned to the norms and preferences of groups and communities. We first examine concepts from the legal realm of case law and note key aspects that we adapt to create CLG. Then, we provide high-level details around the process for general decision-making following the CLG framework. Finally, we detail how CLG could be applied in two different settings: a human-led decision process for communities and an automated process that applies CLG with general-purpose large language models (LLMs).
3.1. Basis in Concepts from Legal Case Law
In judicial systems around the world, high stakes judgments are made regularly that depend on interpreting and applying concepts defined by society. In order to make consistent judgments on cases, some judicial systems have adopted the concept of “case law”, where in addition to higher-level statutory laws, past court decisions on substantially similar cases can also justify a position on a case.
We observe that case law systems rely on two main features to ensure the consistency and legitimacy of judgments without an exhaustive set of laws:
  • “Stand by what has been decided” (or Stare Decisis): Having a consistent criteria is crucial to any dependable system of making judgments—people want similar cases to have similar outcomes in order to perceive the system as fair. Case law systems promote stable judgments by encoding decision criteria implicitly into precedents. Rather than reason about the applicability of abstract criteria, which can vary widely due to subjective interpretations, judges can instead reason about the more constrained question of whether a case is similar to this or that decided case.
  • “Trust but verify” through Legal Citations: Making judgments often involves choosing between opposing options which naturally result in disagreements. When drafting a decision in the legal space, lawyers and judges provide citations to prior cases not just as context but in order to situate the case in the broader graph of dependencies between cases. In this way, the validity of cited cases provides an attestation to the soundness of a decision. By maintaining proper citations, the rationale for decisions can be inspected by anyone through following citation chains.
However, because of the high-stakes nature of cases in the legal system, judgments are typically conducted by experts with years of training and carried out in formal settings where significant time is allotted for deliberation. We must make adaptations to traditional case law before we can apply it to decision-making tasks such as content moderation or alignment of AI models that have significantly higher throughput but less existing structural support (such as detailed reasoning around case judgments).
Figure 2. A diagram showing, at a high level, the process of case law grounding reflected as 3 steps: (1) finding potentially relevant past cases, (2) selecting precedents, and (3) finalizing precedents into judgments. We also show two instantiations of the process: A human-centric process (top) where, after relevant cases are retrieved, human annotators use an interactive tool to view and classify cases, and an AI-centric process (bottom) where an LLM takes the role of a decision-maker and, through chain-of-thought style prompts, evaluate precedents and form decisions.
[…]
A diagram of the case law grounding process implemented both as
3.2. Using Cases to Ground Decision-Making
3.2.1. Step 1: Retrieving Relevant Cases
The first step in case law involves conducting legal research. In the legal practice, lawyers spend considerable effort to locate, examine, and reason about the space of past decided cases in order to draft the best argument for their client. This often involves querying legal databases to identify relevant cases that can be binding precedents for the case they are arguing (Berring, 1987). Once relevant cases are found, there can be additional criteria and checks to make sure that it can apply.
Semantic Similarity for Case Retrieval: Experts in the legal space depend on specialized keywords and indexing systems that let them navigate well reasoned facets of case law to quickly localize their search space. These indices are valuable because they can encode crucial aspects of a case that directly help identify its viability as a precedent (e.g., is the jurisdiction the same, are there commonalities in the circumstances, evidence, or statutory nature of the case); thus it makes sense to spend resources to create them. In the settings we consider, we can’t expect to have ready-made case indices, but we can utilize advances in general purpose language models to approximate case retrieval through text embeddings. Since the cases we’re interested in (e.g., online comments, interpersonal advice scenarios, user queries, and AI model responses) encode all of their factual properties through their natural language descriptions, a reasonable proxy for the relevance of any case to any other can be approximated in the form of the distance between them in the embedding space $d(x_{\text{past}},x_{\text{current}})$. Depending on the domain, it may be possible to choose an embedding specifically created for the domain, though in many situations, an off-the-shelf generic embedding can also be sufficient.
3.2.2. Step 2: Determining Applicable Precedents
Simply surfacing potentially relevant cases to the agent making a decision is likely not sufficient, as we don’t know whether the decision-maker considers the prior case as precedent setting and whether they have integrated it into their judgment. In the legal realm, the selection of precedents from related cases is made explicit, taking the form of a written document containing the arguments being made, and especially the citations to prior precedent cases that support such arguments. Indeed, this explicit identification and documentation of precedents as supporting arguments to a judgment can confer many benefits. For one, by documenting the exact cases used to support a judgment, it becomes easier to audit the reasoning behind judgments. One can gain understanding of disagreements between different decision-makers by inspecting the commonalities and differences in the precedents selected. Additionally, selecting precedents decouples the judgment process from a more subjective (and possibly un-grounded) interpretation of a case and outcome to instead focus on reasoning about more objective comparisons between aspects of the cases. Finally, by having a shared set of precedents to draw from, we can improve consistency of decisions across different adjudicators who may have different personal subjective preferences.
However, requiring a full argument for each case, while effective for challenging crowd judgment tasks (Drapeau et al., 2016; Chen et al., 2019), can be labor intensive. To mitigate this issue, in this stage we simplify the argument process to focus only on classification of whether a retrieved case is a good precedent, not why it is so. This means that instead of coming up with a judgment, writing a justification, and citing cases to support it, an adjudicating agent focuses on presenting their evaluation of the neighborhood of cases.
Specifically, in this step, the agent will be asked to classify the retrieved cases into either:
  • Precedents: A retrieved case should be marked as a precedent if that prior case provides support for the judged case to adopt a similar decision due to substantially similar circumstances or aspects (i.e., cases that fall on the same side of the decision bound of the judged case).
  • Doesn’t Apply: A retrieved case does not apply as a precedent if there are critical aspects of the case that differ from the judged case such that the argument for the past decision would not apply to the new case. While many cases that don’t apply as precedents are likely to fall on the opposite side of the decision bound of the judged case, this is not a requirement for the case to not apply. Sufficiently significant differences in circumstance or factual aspects don’t necessarily need to result in opposing outcomes (e.g., involvement of a minor versus adult in a legal case might affect the argument but the outcome judgment could end up being the same).
While case selection can be led by a human annotator, with advances in reasoning with natural language using language models, it is also possible for an LLM to be used with a specially designed prompt to complete this step.
3.2.3. Step 3: Deriving Judgments from Selected Precedents
Once the sets of precedent cases have been selected in the step above, a final decision can be synthesized by checking each case in the precedent sets and adopting the judgments associated. With the assumption of stare decisis, we view the previous judgments on any retrieved past cases as a reliable source of truth. This means we can view each case as evidencing a direction to take the final judgment.
3.3. Supporting Human-Led and LLM Prompted Decision-Making
(a) Task and instructions
[…]
Screenshot of the task and instructions of the interface, with general instructions on how to use the annotation tool at the top as a bulleted list,
(b) Case law grounding interface for the case condition
(c) Constitutional grounding interface for the rule condition
Figure 3. Screenshots of the annotation tool interface used to collect annotations in the human-led process.
3.3.1. Case Selection Tool for Human-Led Decision-Making
For the human-led decision-making process, we built an annotation interface that is deployed to our participants in the form of a web application (Figure 3). We split the annotation interface down the middle, presenting the general usage instructions and the case being judged on the left (3(a)), and interface components to support making grounded decisions placed on the right.
We use the right side interface to facilitate the steps for conducting decision-making following or case law grounding process. During annotation, the right side of the interface is populated with a list of cases retrieved through the case retrieval step (3(b)). Under each retrieved case, the annotator can click on a button to toggle whether a case should be considered a precedent or not. Once the annotator is done evaluating each case, they are asked to provide a final decision given their review of the precedents but not constrained to the decisions of them. While selecting the set of applicable precedents alone is sufficient for the CLG workflow, this last question is included to allow us to evaluate whether enforcing binding precedents (subsubsection 3.2.3) is important to final accuracy or not.
Additionally, for our experiments, we adapted the same interface to also facilitate decision-making following a more traditional constitutional grounding process. In the constitutional grounding condition, we use the same general interface but render a list of rules and guidelines associated with each case in our task domain (3(c)) instead of retrieved precedents. We also include checkboxes for each rule, prompting the human annotator to indicate any that they applied during their decision-making process. As we can’t derive the final decision directly in constitutional grounding, we rely on the annotators to provide their own final decision, collecting this decision using the same design as the case law grounding version.
3.3.2. LLM Prompts for Precedent Selection
To evaluate an automated process where LLMs contribute to decision-making instead of human annotators, we also implemented a series of prompts for LLMs to follow a grounding process similar to the human-led version (Appendix A). In the prompting version of the workflow, we set up the system prompt in a similar way to the instructions section in the annotation interface, providing the instructions for the decision task itself (whether the model is assessing case precedents or constitutional guidelines) as well as task domain specific instructions. We then provide the case being judged along with the grounding support as part of the user prompt. Each user prompt assessment was prompted independently, with no conversation history persisted between two different user prompt queries.
For the LLM-prompting version of CLG, we use a case selection prompt (subsection A.2) to evaluate each retrieved case, substituting {input} with the case being judged, {precedent} with one retrieved past case, and {decision} with its past decision. This prompt was then repeatedly invoked for each retrieved case, producing a binary output for whether the case should be considered a precedent or not. Following the CLG process, these precedent determinations were then used to derive the final decision. Unlike the human annotation interface where annotators also produced a final decision based on their review of precedents but unconstrained by the decisions of those precedents, we could not create a prompt that provided all precedents to the model to produce a single final decision in a way that fit within the context window of the model, thus non-binding precedents were only evaluated for the human-led process and not the LLM prompting version. For the LLM-prompting version of constitutional grounding, we use a guideline assessment prompt (subsection A.3) to evaluate each retrieved case, substituting {input} with the case being judged and {instructions} with a list of rules and guidelines for making the decision. We then collected the resulting output directly as the model’s final decision.
4. Experiments
As the goal of CLG is to support decision-making on subjective and socially-situated tasks for groups and communities, we focused our evaluation of CLG on these types of tasks, drawing from two common subjective decision tasks: content moderation and toxicity rating. We then focused on sampling several distinct groups or communities which have been observed by prior work to have distinct preferences and norms when it comes to making decisions on these subjective tasks to compare decisions made with support from traditional constitutional grounding (rule) against decisions made with case law grounding (case) support. We set out to answer the following research questions in settings where humans are the decision-makers as well as settings where LLMs are being prompted to make decisions:
  • RQ1: For subjective and socially-situated decisions, does the case law grounding approach result in more accurately aligned decisions compared to traditional rule-based constitutional grounding?
  • RQ2: For case law grounding, how do different choices in the retrieval setup and precedent selection affect the alignment accuracy of decisions produced?
    • RQ2a: How does the alignment accuracy scale with different retrieval window size ($k$) for precedents?
    • RQ2b: How does enforcing binding precedents affect the alignment accuracy of the decisions?
  • RQ3: For the automated LLM-prompting based process, do our findings around CLG continue to scale as we move to larger datasets / additional population groups?
4.1. Experiment Setup
For RQ1, we set up our experiments to focus on evaluating how well the decisions produced by different grounding approaches aligned with the ground truth, comparing two main conditions rule—which represents a constitutional grounding process, and case—which represents the case law grounding approach in CLG. To more comprehensively evaluate these two conditions, we conducted this comparison across 5 groups or communities each of which were making decisions on one of 2 task domains. In addition to groups, we also evaluated whether our findings between conditions holds across different versions of the process that used different agents for making the decisions: human-led and llm-prompting based. For RQ2, we examine how different configurations for CLG may affect performance, with RQ2a assessing different retrieval window sizes via simulation and RQ2b comparing an additional condition example for the human-led process that represents a weaker form of case law grounding. Finally, for RQ3, we only look at the LLM-prompting version of both CLG and constitutional grounding, comparing whether they retain their relative performance on a additional sets of cases that we did not collect human annotations for.
4.1.1. Tasks and Datasets
We picked two real-world decision-making tasks that involve subjective decisions with social context specific to groups: a comment moderation task where the decision is to “keep” or “remove” a comment posted in a Reddit community, and a toxicity rating tasks where the decision is the scalar rating of the toxicity level of a social media post (on a scale from 1–5). Then, for each task, we selected a dataset that contained ground truth annotations provided by: (1) different groups of users, (2) that we could recruit for our annotation study.
For the mod domain, we used the NormVio dataset (Park et al., 2021) which contains comments from various communities on Reddit that may or may not have violated the norms of that community and as a result have been evaluated by moderators for potential removal. This dataset separates comments by the “subreddit” it was posted in, which we use to partition distinct communities that will likely have distinct preferences around content moderation. From this dataset, we selected the top 10 subreddits with a sufficiently large size of examples ($n>100$) and did not exhibit significant skew towards either decisions of “keep” or “remove”. We then further narrowed down this set based on avoiding multiple communities around similar topics and prioritizing communities that were easier to recruit from, ultimately sub-sampling this dataset down to 3 communities: r/Coronavirus, r/AmItheAsshole, r/Gaming. Ground truth labels for each case is established by the actual moderation decision applied to each comment.
For the tox domain, we used the toxicity perspectives dataset (Kumar et al., 2021) which contains a wide range of posts collected from social media platforms that were then evaluated by crowd annotators and rated for toxicity. This dataset contains disaggregated ratings that are associated with demographic information of the annotator, which we use to segment a general crowd worker population into distinct salient groups. Based on the findings from the original paper, we sampled from 3 disjoint intersectional demographic slices (see subsubsection 4.1.2) that were indicated to have the most significant differences with each other: seen=N,targeted=N, seen=Y,targeted=N, seen=Y,targeted=Y. Ground truth labels for each case in each group is established by taking the mean rating for that group rounded to the nearest integer.
For each slice in each task domain above, we first shuffled the instances randomly, then drew 10% of cases only used as past precedents, 10% judged by both the human-led process and LLM prompting (rounded down to nearest multiple of 10 to facilitate annotation), and 10% only evaluated with LLM prompting. To ensure a consistent level of effort for the annotation task, cases set to be judged by human annotators were partitioned into batches of 10, with each annotator only judging one such batch within their session. When evaluating the constitutional grounding condition (rule), we use published community guidelines for each subreddit as approximations of a high-quality constitution for groups in mod. Meanwhile, for the tox domain, we adapt the specific instructions and criteria given to annotators for rating toxicity in the original dataset from Kumar et. al. as a list of rules that serve as the constitution for all groups. Finally, to address the fact that community guidelines may have changed over time, we spot checked the community guidelines for each of the 3 groups in the mod domain using the Wayback Machine [^1] https://web.archive.org/ to confirm that no significant changes (such as adding or removing rules) were made to the guidelines across the time span of cases evaluated.
4.1.2. Communities and Groups
One challenge of decision-making around socially-constructed criteria so challenging is the abundance of localized preferences over criteria that differ across different population groups or communities. As CLG is intended to support making decisions around subjective concepts in a way that aligns with the norms and preferences of a group, we set up our evaluation to include two types of social groupings that are likely to have consistent norms and preferences around subjective decisions: interest-based communities—consisting of participants recruited from active users within a subreddit, and demographic-based population groups—consisting of general participants that report belonging to some demographic group.
For the task of content moderation (mod), where decisions are made around whether content is removed or retained, prior work has observed that different communities can often have significantly different preferences when it comes to adhering to a formally defined criteria or values (Koshy et al., 2023). Thus, drawing from our dataset, we selected 3 communities from Reddit: r/Games—a broad interest group around game content and the games industry, r/Coronavirus—a community focused around high-quality science-informed discussions related to COVID-19, and r/AmItheAsshole—a collective space for sharing and opining over a variety of interpersonal conflict scenarios.
For the task of toxicity rating (toxicity), prior work has found that differences in the background of raters, such as their prior exposure to toxic content as observers or victims, can affect how they rate the same piece of text (Kumar et al., 2021). While raters do not necessarily form a “community”, we can still view demographic features as a way to separate the population into groups that represent distinct norms and preferences. In our case, we stratify the general population by taking two demographic properties identified to be significant correlating factors (Kumar et al., 2021) of toxicity ratings: whether a rater has often seen toxic content when interacting with social media (negatively correlated with scores), and whether the rater has been personally targeted by toxic content (positively correlated with scores). Noting that it does not make sense to have been targeted yet never seen toxic content, this results in 3 final groups: seen=Y,targeted=N, seen=N,targeted=N, and seen=Y,targeted=Y. During recruitment, we found it challenging to find participants for seen=N,targeted=N, so we only evaluated the LLM-prompting process (RQ3) with that group, excluding it from other evaluations.
4.1.3. Participant Recruitment and Crowdsourced Annotation
As we envision human-led processes for grounding subjective decisions (whether constitutional or case law) to be used in practice by groups and communities themselves rather than by external populations, we set out to recruit participants that would demographically best match the original group or community producing the ground truth data. To achieve this goal, we applied pre-screening surveys in each of our recruitment efforts (Appendix B), with participants first asked to fill out the screening survey, after which we would select and assign tasks conditions to them.
As each of the mod groups represents a specific subreddit community, we recruited participants by running Reddit Ads targeted to subreddits used in our task domains. Participants were asked to fill out the screening survey to self report subreddits they frequented within a list of communities (subset of the original top-10 candidates subsubsection 4.1.1, removing 1 subreddit that had been removed). Then, participants who reported frequenting at least one of the 3 target subreddits would be allocated a batch and condition, balancing such that each community, condition, and batch had a target of 3 annotators. At the end, this resulted in annotations on 3 batches for r/AmItheAsshole, 2 for r/Games, and 1 for r/Coronavirus.
In the tox domain, original annotations were provided by Amazon Mechanical Turk (AMT) workers. However, recent work has shown concerns around AMT (Marshall et al., 2023) annotations that have developed since the original dataset was collected. To avoid issues with introducing significantly more quality control components than the original task, we opted to recruit from Prolific, as it is a similar crowd work platform with (as of our deployment) lower rates of data quality issues. Participants were screened with two questions about their prior exposure to toxic posts with these answers then used to assign them to one of 3 population groups (see 4.1.2). During recruitment, we also found that no crowd workers reported being in the seen=N,targeted=N group during screening. While this was ultimately unsurprising, we opted to drop this group from the evaluations related to human-led processes (RQ1 and RQ2), and only used this group for the final LLM experiments around scaling.
As both experiment domains involve reading content that is potentially offensive, we made sure to prominently disclose this aspect during recruitment. We also allowed participants to drop out during the task, if they experienced discomfort with the annotation. Human participants in both conditions were paid a base rate of \$15 for completing our annotation study (30-45 minutes), with Prolific participants being paid through the platform and Redditors being paid in the form of an Amazon Gift Card. Annotators electing to withdraw consent during the task were pro-rated payment for the amount of annotations completed, with their data excluded from the evaluation. Additionally, to incentivize participation from Redditors and reduce drop-out between screening and completion, we provided an additional bonus of \$5 to Redditors who completed our annotation task within 24 hours of assignment. In total, we were able to recruit $n=60$ participants from Prolific and $n=35$ participants from Reddit.
4.1.4. Additional Setup Details
Language Model and Embedding Model. As our goal was not to evaluate a variety of language models, we opted to keep the models used to produce embeddings and the LLM used to execute prompts consistent and fixed across all conditions and task domains. For the embedding model, we used the general purpose embedding offered by OpenAI [^2] https://openai.com/blog/new-embedding-models-and-api-updates (“text-embedding-3-large” ) with a distance metric based on cosine similarity for ranking relevant cases. As precedent retrieval only depends on the case being judged, we pre-computed the ordered list of retrieved past cases with $k=15$ as the retrieval window and provided this for human annotations, with smaller window sizes simulated post-hoc. For the LLM used during prompting, we used the gpt-4-turbo-preview model offered by OpenAI.
Retrieval Window Simulations. In RQ2a, we evaluate how alignment accuracy may be affected by the retrieval window size, specifically examining different options of $k\in\{1,3,5,10,15\}$. However, directly evaluating different retrieval outputs in the human-led workflow would be very costly, especially given the observation that the set of retrieved examples at smaller $k$ values would just be subsets of those at larger $k$. Thus, instead of separately testing different values of $k$, we only have the human participants and LLM select precedents over the largest retrieval window size $k=15$. We then simulate the outcome of precedent selection for smaller $k$ by dropping any selected precedents that would fall outside the retrieval window at that size.
Consensus Tie-Breaking for Binding Precedents. When deriving final decisions based on precedents selected, it is possible for the collection of precedents selected by either humans or LLMs to contain a plurality of decision outcomes. While in the legal realm such conflicts would indicate further reasoning and deliberation, in CLG we apply a simpler deterministic approach to resolve conflicts based on approval voting. We collect the past decision of each selected precedent, and then take the mode (most common value) to be the final decision. In the event of a tie where there are multiple values for the mode, we break the tie by selecting the decision that has a corresponding precedent case that is closest (retrieval distance) to the original example being judged. Should the set of precedents be empty, we treat this case as a tie between all retrieved cases with each getting no votes.
4.2. RQ1: Comparing Case Law and Constitutional Grounding
(a) Accuracy of human decisions the tox domain
(b) Accuracy of human decisions the mod domain
Figure 4. Accuracy of the decisions produced by humans using each grounding approach, rule for constitutional grounding and case for Case Law grounding. For the case condition, results reflect the retrieval setup used during annotation ($k=15$). We also include an example condition, which presents the accuracy when binding precedents are not enforced and retrieved cases primarily act as examples (details discussed in subsubsection 4.3.2).
[…]
Bar plots comparing the accuracy of decisions produced by humans using each grounding approach.
(a) Accuracy of LLM decisions the tox domain
(b) Accuracy of LLM decisions the mod domain
Figure 5. Accuracy of the decisions produced by LLMs prompted using each grounding approach, rule for constitutional grounding and case for Case Law grounding. To match human decision-makers, $k=15$ was also used as the retrieval window for presenting precedents to the LLM in the case condition.
[…]
Bar plots comparing the accuracy of decisions produced by LLMs using each grounding approach.
To answer RQ1, we compare the accuracy at which each grounding approach reproduces the ground truth decisions observed for each group and task. For the human-led process, we follow the workflow in subsection 3.3, using human annotators for each task and group to produce decisions grounded by either constitutional grounding (rule) or case law grounding (case). We found that for the human-led process (Figure 4), case law grounding was able to produce statistically significantly higher accuracy in 4 out of 5 groups: +16.7 p.p.333percentage points for seen=Y,targeted=N in tox (paired t-test, $p=0.010<0.05$), +19.1 p.p. for seen=Y,targeted=Y in tox (paired t-test, $p=0.019<0.05$), +23.3 p.p. for Coronavirus in mod (paired t-test, $p=0.045<0.05$), and +16.0 p.p. for Games in mod (paired t-test, $p=0.029<0.05$). However, we did not observe any significant difference between the two grounding approaches in the AmItheAsshole in mod, with a difference of 3.3 p.p. (paired t-test, $p=0.721>0.05$).
For the LLM-prompting process, we also follow the prompts in subsection 3.3. We found that for the LLM process (Figure 5), case law grounding was able to produce statistically significantly higher accuracy in 4 out of 5 groups: +23.6 p.p. for seen=Y,targeted=N in tox (paired t-test, $p=0.004<0.05$), +20.8 p.p. for seen=Y,targeted=Y in tox (paired t-test, $p=0.006<0.05$), +23.9 p.p. for Coronavirus in mod (paired t-test, $p=0.011<0.05$), and +32.9 p.p. for Games in mod (paired t-test, $p=0.006<0.05$). However, similar to the human-led version, we did not observe any significant difference between the two grounding approaches in the AmItheAsshole in mod, with a difference of +10.6 p.p. (paired t-test, $p=0.229>0.05$).
Overall, these results suggest that for the vast majority of the tested groups and tasks, case law grounding did result in decisions that had statistically significantly higher accuracy compared to the ground truth. We also find AmItheAsshole in mod to be an outlier—while average accuracy was still higher the difference was not significant, which indicates that there may be communities and tasks where cases are not as effective. Coupled with results from RQ2b, we discuss in subsection 5.1 how limitations around the case law grounding process, such as challenges in situations where individual cases are longer and more complex, may have contribute to the reduced effectiveness of using cases as the main grounding mechanism for groups like AmItheAsshole.
4.3. RQ2: Assessing Retrieval and Decision Derivation Configurations
For our second RQ, we examine two configuration aspects of the CLG workflow that may contribute to the final accuracy: the retrieval window size ($k$), and whether to enforce binding precedents (case) as is default in CLG or not (example).
4.3.1. RQ2a: Alignment Accuracy and Retrieval Window Size
(a) Accuracy scaling on the tox domain for humans
(b) Accuracy scaling on the mod domain for humans
Figure 6. Scaling of human decision accuracy given different retrieval window sizes $k\in\{1,3,5,10,15\}$ for Case Law grounding (simulated results when $k\neq 15$). We also plot two reference lines: oracle indicates the upper-bound accuracy when picking the best available precedent(s) within the retrieval window; knn indicates the accuracy of the consensus decision for each retrieval window. Additionally, we note that rule does not involve retrieval and thus has no window size ($k$), however we plot it at $k=15$ to match the non-simulated window size $k$ when collecting results for case.
[…]
Line plots
(a) Accuracy scaling on the tox domain for LLMs
(b) Accuracy scaling on the mod domain for LLMs
Figure 7. Scaling of LLM decision accuracy given different retrieval window sizes $k\in\{1,3,5,10,15\}$ for Case Law grounding. Reference lines for oracle and knn same as in Figure 6. rule indicates outputs for LLM prompts with constitutional grounding, also plotted at $k=15$.
[…]
Line plots
In Figure 6 and Figure 7, we plot the simulated accuracy of CLG (case) in the human-led and LLM-prompting based processes at smaller retrieval window sizes than we used for RQ1 to evaluate whether smaller retrieval windows may be sufficient. In addition to plotting the scaling of case at different values of $k$ and presenting the rule condition for comparison at $k=15$, we also plot how 2 additional reference lines behave at different values of $k$: oracle and knn. As CLG uses binding precedents, the oracle reference line indicates an upper bound on accuracy given a certain retrieval window. For any case being judged, if the retrieval window contains any past case with the ground truth decision, oracle will always produce the correct answer, otherwise it will always be incorrect. We can observe that at small retrieval windows, it is often the case that across the entire set of judged cases, even the oracle cannot achieve perfect accuracy, as for some cases no correct selection of precedents exists. However, as $k$ becomes larger, we see oracle more consistently achieving perfect accuracy. The knn reference line represents a strategy that always selects all precedents and relies on the tie-breaking mechanism in CLG to produce the final decision. Deviations between case and knn indicates the effect of contributions from the decision-maker as a result of precedent selection.
With both human and LLM processes, we did not find any consistent correlation between $k$ values and accuracy across the tested groups and domains, nor did we find statistically significant differences between accuracy of neighboring $k$ values in the set of values tested. This suggests that for many of the individual groups and domains tested, smaller $k$ values may be sufficient to achieve a similar level of accuracy. However, looking across groups we also observed the following behavior patterns that we believe suggest against picking very small $k$ values to reduce decision cost, especially since it is not possible to evaluate scaling in practice when ground truth is not available:
Performance plateau behavior. In most of the groups and domains (excluding the outlier AmItheAsshole we observed in RQ1 evaluations), accuracy tends to eventually reach a plateau where further scaling of $k$ contributes little to accuracy. However, the speed at which this is reached varies by the group and domain: For example, groups like seen=Y,targeted=N or Coronavirus seem to reach a plateau at $k=3$ and $k=5$, while Games still sees accuracy improvements at $k=10$. While this behavior not too surprising by itself—with more dissimilar cases retrieved, we should expect additional cases to be less and less relevant to the final decision—there does not seem to be a consistent $k$ for reaching the plateau across groups and tasks. We hypothesize that this variation across groups is likely to do with aspects like the “density” (average distance) of past cases that can be retrieved and distributional properties of past cases around the judged cases.
Unreliable for retrieval windows that are too small. An additional observation is that, at very small values of $k$ such as $k=1$ or $k=3$, performance can be relatively unpredictable. In some groups, like seen=Y,targeted=Y and Coronavirus, we see initial drops in performance at lower $k$ values. While this drop is not significant when viewed across all the judged cases, it does indicate that accuracy is much more sensitive to $k$ when $k$ is small—moving between retrieval thresholds can impact accuracy on some portions of cases. We hypothesize that this may be due to case-based approaches naturally being more sensitive when the retrieval window is low, as small changes could result in relevant precedents not making it past the threshold or immediate local neighborhoods primarily comprising of similar but less relevant cases.
We discuss the implications of some of these observations in subsection 5.1, noting potential areas of future work that may provide signals for optimal $k$ selection without relying on an annotation-intensive process like how we are evaluating RQ2a, where an examination of scaling is conducted using a subset of cases where ground truth is known. Finally, we also found that the AmItheAsshole group (which was an outlier before) still to exhibits distinctive behavior—we see accuracy first improve as the window size increases but then fall back down (though still not to a statistically significant degree, with even the top performing $k$ still not being statistically significantly higher performance than rule, $p=0.079>0.05$). However, as this behavior happens at small $k$ ($k=3$ and $5$), we believe this is reflects the earlier observation of unpredictable performance when $k$ is small.
4.3.2. RQ2b: Binding Precedents versus Cases as Examples
In prior work around grounding human decision-making in crowdsourcing settings, examples (Wu and Quinn, 2017) have been noted as an important feature for constructing more effective guidelines or “constitutions”. Unlike CLG, however, constitutional grounding examples are selected and incorporated during the construction of the constitution, meaning that they do not adapt to properties of the case being judged. This raises the question of whether simply retrieving and presenting prior related cases for a new case in the form of examples alone is sufficient for improving grounding accuracy or if the binding nature of precedents—where decisions must derive directly from selected precedent sets—needs to be applied for CLG to produce accurately grounded decisions.
To examine this, during the human annotation process, we not only collected participants’ judgments on whether each precedent applied, we also asked participants to give a final decision on the judged case after reviewing the list of retrieved prior cases without being constrained to the consensus of the the cases they selected as precedents (as noted in subsection 3.3). In Figure 4, we include an additional condition—example—to reflect the accuracy of this configuration. Comparing the accuracy produced by non-binding precedents (example) to that of constitutional grounding (rule), we note the absence of statistically significant differences (paired t-test, $p=0.635,0.360,0.458,0.266>0.05$ for seen=Y,targeted=N, seen=Y,targeted=Y, Coronavirus, Games) in most of the groups, despite the fact that we did observe statistically significant improvements to accuracy when using binding precedents (case). Here, again, AmItheAsshole is an outlier, with example showing a statistically significant improvement over rule (+20.0 p.p, $p=0.020<0.05$) even though binding precedents (case) did not show significant improvements.
4.4. RQ3: Assessing the LLM Workflow with Larger Datasets
(a) LLM accuracy on the additional cases in tox
(b) LLM accuracy on the additional cases in mod
Figure 8. LLM performance and scaling on additional cases for the tox and mod domains. Reference lines for oracle and knn are produced in the same way as Figure 6. rule also reflects LLM performance using constitutional grounding on these cases. Retrieval still done at $k=15$ with lower $k$ windows simulated post-hoc.
[…]
Line plots
Finally, in our last evaluation, we use a larger set of cases (that we did not evaluate the human-led process on) to evaluate the LLM-prompting based process for CLG, examining whether our observations from RQ1 and RQ2 still hold (Figure 8). At the same $k=15$ retrieval window, we find that across all 6 groups at this larger scale, case law grounding (case) consistently outperformed constitutional grounding (rule) when using the prompting version: For the tox domain, we observed +20.7 p.p. for seen=Y,targeted=N (paired t-test, $p=0.0002<0.05$), +24.8 p.p. for seen=Y,targeted=N (paired t-test, $p<10^{-4}<0.05$), and +13.6 p.p. for seen=N,targeted=N (paired t-test, $p=0.016<0.05$). For the mod domain, we observed +9.2 p.p. for Coronavirus (paired t-test, $p=0.0418<0.05$), +13.3 p.p. for AmItheAsshole (paired t-test, $p=0.0047<0.05$), and +30.0 p.p. for Games (paired t-test, $p<10^{-5}<0.05$). We also observe similar trends (plateaus and more variations at low $k$) with respect to scaling over $k$ as we did for RQ2a, with the only major difference being that with this larger dataset, we do now see improvements sufficient for statistical significance even with the AmItheAsshole that was previously an outlier without significant results.
5. Discussion
In this section, we will discuss some of the limitations of case law grounding and how they may have contributed to some of our observations in our experiment results. Then, we will discuss implications and considerations around the human-led decision process versus LLM prompting. Finally, we will discuss how CLG and constitutional grounding can be complementary solutions to the problem of grounding subjective decisions to group or community norms and preferences, and how we can view the two processes within the framework of policy-making in communities.
5.1. Limitations of Cases as Grounding Mechanisms
As we found in our experiments, while CLG was effective compared to constitutional grounding in most of the group decisions evaluated, there were also situations where CLG did not perform significantly better than constitutional grounding. One possible explanation for this may lie in limitations of cases themselves when it comes to grounding decisions. While the concrete nature of cases means it may be less ambiguous for decision-makers (human or LLM) to assess each case, the specificity encoded in cases can also have several drawbacks.
For one, cases contain details and context that support making the decision, meaning they can be longer and more complex. Since CLG uses similarity-based metrics for retrieval, longer cases means higher chances of spurious similarities being picked up by the retrieval process, degrading the quality of set of retrieved cases. Compared to other groups in mod, AmItheAsshole is a community around interpersonal conflicts, and as a result some comments contain much more descriptive details while others do not. Additionally, because these comments contain complex context, it is likely that even when a precedent is relevant, it may not be associated with the right “keep”/“remove” decision due to other details of the case. Considering the observation that non-binding precedents performed significantly better while binding ones did not, this limitation may have been an important factor.
Furthermore, CLG performance can depend on the (local) density of past decisions around each new decision in the retrieval window. As we observe in RQ2a, CLG can perform well for some groups even with small retrieval windows, thus saving on costs (in human effort or model computation). However, window sizes that are too small risk missing relevant precedents while too big of a retrieval window risks introducing irrelevant and confusing prior decisions. While our evaluation setup in RQ2a conveniently allows us to evaluate different values of $k$, in a practical deployment we would not have access to the ground truth decisions and thus may end up with a sub-optimal $k$.
Finally, beyond mechanistic limitations of cases, using case law as the main grounding mechanism can also incur social or emotional burdens on human decision-makers. In the domain of the content moderation, prior work (Steiger et al., 2021) has noted that simply assessing harmful and disturbing content imparts an emotional toll on moderators. With case law grounding approaches, we may end up exposing human decision-makers to a higher amount of harmful content during their precedent selection compared to constitutional grounding. While we took measures to remove identifiable information in our experiment, some annotators still expressed concerns around privacy when showing full text of past posts or comments.
We believe there are many opportunities for future work to address these limitations, ranging from using new metrics to estimate optimal $k$ to LLM-based tools rewrite cases while balancing removal of irrelevant or identifiable context.
5.2. Large Language Models for Community Decision-Making
In our experiments, we observed that despite not being specifically prompted or tuned for a particular group or community, the LLM version of CLG (case) still performed relatively well, in certain cases slightly outperforming human annotators when it comes to average accuracy (though overall we did not observe statistically significant different accuracy between CLG with human decision-makers or CLG with LLM prompting, with both performing on-par with each other). Looking at the groups that saw the most difference between human and LLM performance, we note a couple of potentially interesting differences:
In AmItheAsshole, the LLM process produced slightly higher accuracy for the CLG (case) condition. For many of the cases where the LLM produced a correct decision while the human decision-maker did not, the set of retrieved past cases often contained more complex and longer cases, coupled with the fact that a $k=15$ retrieval window was used during annotation, this possibly suggests that for CLG with more complex cases, human decision-makers may be making more-than-expected errors due to possible fatigue. In Games, we also observe that LLMs did much worse with constitutional grounding (rule) than human decision makers. Examining the constitutional guidelines, we found that, unlike other groups, Games had a very long and complex set of community guidelines (spanning a total of 27 individual rules with detailed descriptions for each as opposed to around 10 rules for Coronavirus and AmItheAsshole). We hypothesize that in this case, human decision-makers were better able to intuit many of the complex rules quickly, whereas the much longer and detailed set of instructions may have made it harder for LLMs to apply the correct criteria.
Overall, while we did not find any clear patterns that suggest LLM prompting makes more specific types of failures (at a group level) compared to human decision-makers when using CLG, we do notice that human decision-makers did slightly better when using constitutional grounding than LLMs. We hypothesize that this may be due to a combination of the fact that human participants were drawn from members of the relevant groups / communities, and thus may have internalized some of the norms and preferences of the group, coupled with more familiarity when consuming community guidelines which were written with human understanding in mind. Given these observations, we believe that if groups and communities wish to start adopting LLMs to automate decision-making, following more of a case law approach rather than supplying existing constitutions made for humans to a LLM is likely to result in better results. For the groups and communities we evaluated, past decisions seem to be a much stronger signal for grounding future decisions than their constitutional guidelines, though this may not generalize to more marginalized groups or communities around more sensitive topics.
5.3. The Complementarity of Constitutions and Case Law
In this work, we explored how a case law inspired approach to grounding the decision-making process in socially defined contexts can lead to closer alignment towards group and community preferences when compared to a purely constitutional approach using existing guidelines. However, constitutions and case law do not oppose each other in practical applications, and in legal systems it is not uncommon to have mixed systems that utilize both in conjunction (Palmer, 2012). Thus when looking at decision-making within socio-techincal systems, it is also useful to see both constitutions and case law as complementary means for grounding decision-making, with their own strengths and weaknesses.
Take the example of content moderation in online communities: It’s common to see communities organically utilize constitutional grounding in the form of “community guidelines”. However, others have also observed that community guidelines are not always enforced consistently by moderators in the wild, and not all community members agree with all the specific norms and criteria described by the community guidelines (Jhaver et al., 2019a). Nonetheless, community guidelines are still a helpful means for communities to convey values and norms to both existing members and newcomers alike, even if they may not have the comprehensiveness required for making certain situational decisions. Here is where we believe case law and constitutions can act as complements to each other: The concise abstractions offered by constitutions help ground general values and preferences when making decisions, while the specificity and nuance offered by precedents help clarify how to consistently apply these values in-context. In fact, during our collection of community guidelines on Reddit for our study, we also observed communities including community-specific examples within their constitutional guidelines as a tool for clarifying concepts or making situational exceptions—both aspects we believe a case-based approach like CLG excel at.
Having constitutions and case law as complementary components can also be useful in situations that demand socially-appropriate decisions without complete human supervision. The recent commercial deployments of large language models (LLMs) as general purpose AI agents has also brought along high-profile cases of models failing to consider the social appropriateness of responses (Weidinger et al., 2021). Unlike with content moderation, it is impractical to have human decision-makers determine appropriateness for each user query and response as they come. Accordingly, providers of these AI agents have been working on ways to ensure that the outputs of such systems also follow the norms and expectations of broader society based on constitutional approaches as well. One such example of this is with Constitutional AI (Bai et al., 2022), where the general public is tasked with helping define universal values into a constitution that is then incorporated into model training and tuning. Here again, the case-based approach in CLG could provide additional utility where the LLM-backed decision workflow would allow both constitutions and retrieved in-context examples to be applied during inference time to determine appropriate responses or strategies for responding, in a way similar to recent work with retrieval augmentation for generative models and in-context learning (Lewis et al., 2020b; Lin et al., 2024). This could help bridge potential issues where models may “interpret” human-produced guidelines differently, by also introducing relevant human-made decisions.
5.4. Addressing Consensus and Evolution of Preferences
In our current implementation of CLG, we use a simple approval voting mechanism to resolve conflicts around the final decision given the precedent set. However, in many situations conflicts can be a signal that past cases may be insufficient to ground the current decision and that we must clarify the decision bounds. While CLG currently does not include a mechanism for this kind of consensus-building, we can look towards prior work for possible implementations that could be incorporated into CLG. For example, when conflicts demand clarity on new decision criteria, voting systems like digital juries (Pan et al., 2022) could be used to engage opinions of different stakeholders. On the other hand, when conflicts arise from disagreements when reasoning about how to apply past cases, deliberation may be the tool to resolve them (Chen et al., 2019).
Another common criticism of case law in general centers around issues with the doctrine of stare decisis (Douglas, 1949; Kozel, 2010). Since societal and community values tend to evolve over time (Lin et al., 2017), if future decisions must always be bound to past decisions, then case law decisions would grow increasingly detached from current values. To address this in the long term, case law grounding also needs ways to retire or overturn past precedents. Some have even pointed to this flexibility as a benefit of case law when it comes to quickly adapting to shifting preferences of the public (Cardozo, 2011).
With relatively minor modifications, the CLG process could be adapted to support decision-making over a long period of time by automatically retiring old cases and decisions that have become “stale” and rarely cited and overturning past decisions through a simple escalation process during precedent selection—in addition to indicating relevance, decision-makers could also flag precedents whose decisions they disagree with. By additionally including book-keeping around precedent dependency chains, retired or overturned cases can also automatically trigger re-evaluations of past decisions that depended on them.
5.5. Policy-Making through Case Deliberation
One aspect that differentiates of constitutional grounding from case law is how constitutions often facilitate policy-making (Zhang et al., 2020) in addition to being a way to ground decisions on policy. Groups and communities often use constitutions and guidelines as a way to formalize their norms and preferences, with the process of making constitutions a potentially important component to reflect on the values of the group or community (Butler et al., 2008). In these respects, constitutions can have the advantage of being concise, and thus easy to form opinions on, as well as abstract, allowing coverage of a large part of the decision space with relatively low effort. For larger communities, like that of Wikipedia, there can be complex documented procedures that can be triggered by members to evaluate proposed policy against current practice and used to formalize core policy values into concise, applicable rules or even implemented as bots and algorithms (Hwang and Shaw, 2022). Indeed, the misalignment observed between decisions and guidelines could in part be attributed to the sometimes aspirational nature of constitutions to reflect ideals.
In contrast, while case law can be easy to bootstrap initially—in many cases like content moderation, past decisions are readily available to be used as precedents—building a precedent-based policy solely from naturally encountered decisions can be inefficient, as each precedent only addresses a small part of the decision space. Precedents themselves also offer limited insight into higher-level patterns like community or group values, as each decisions on cases are specific and contextual in nature. However, cases can still serve as a useful medium for policy-making in a different way, by helping groups and communities better consider nuances in decision bounds and “edge cases”. When policy-making only focuses on high-level constitutions, decision-makers are left to interpret ambiguities themselves when policies need to be applied. Because cases are specific in nature, if we can identify or even synthesize cases that lie within the ambiguous spaces between high-level constitutional guidelines, we could surface them during community policy-making processes, where they could contribute to more clarity for the resulting constitution, or, when combined with democratic processes, become binding precedents for controversial situations. We envision complete future applications of case law grounding in community and group settings to also involve case-based policy-making components. Specifically, a case-based policy-making process could analyze existing decisions and precedents to identify areas where policy ambiguity is high or where disagreements are common. These results could serve as signals for selecting or synthesizing challenging cases to collect community inputs on, with the end goal of producing both precedents that directly define decision bounds and potential intermediate probes for building and refining constitutional guidelines.
6. Conclusion
In this paper, we present case law grounding (CLG), a novel workflow that uses precedents for aligning decisions to socially constructed norms and preferences of groups and communities. We evaluate two versions of CLG, a human-led process and an automated LLM process, on 2 decision-making tasks with decisions made by 5 groups. We found that for 4 / 5 groups, CLG performed significantly better than traditional constitutional grounding in accurately producing group decisions. We also examine the impacts of different configurations of CLG such as retrieval window size and binding or non-binding precedents. Finally, we discuss limitations of CLG and how it fits into the broader ecosystem around supporting aligned decision-making in social contexts.
Appendix A LLM Prompts
A.1. Task System Prompts
System prompt for mod case:
You are an AI assistant that is assisting a moderator analyze
comments posted in an online community.
For each comment you are analyzing, you will be shown past
comments that the moderator has either removed or kept.
Your task is to decide whether the decision on each of the
past comments would be relevant to the new comment.
A relevant decision is a decision where the new comment
exhibits some similarities to the past comment.
Think step by step.
System prompt for mod rule:
You are an AI assistant that is assisting a moderator analyze
comments posted in an online community.
The moderator will provide a list of rules that posts and
comments in the community must adhere to.
Based on the list of rules, you should analyze the comment,
and respond with either "keep" or "remove".
Think step by step.
System prompt for toxicity case:
You are an AI assistant that is assisting a user in organizing
their online feed by rating whether content is toxic or not.
For each post in the feed, you will be shown past posts that
the user has rated toxicity for on a scale from 1 to 5.
Your task is to decide whether the toxicity rating of the
past post would be relevant to the new post.
A relevant rating is a rating where the new post exhibits
some similarities to the past post.
Think step by step.
System prompt for toxicity rule:
You are an AI assistant that is assisting a user in
organizing their online feed by identifying whether content
is toxic or not.
The user will provide a list of indicators of what content is
commonly toxic.
Based on the list, you should analyze the content, and
respond with a number between 1 and 5, with 1 indicating
the content is not toxic at all, and 5 indicating that the
content is extremely toxic.
Think step by step.
A.2. Prompts for Case Law Grounding
Prompt for case selection:
The comment you are analyzing is:
{input}
The moderator has judged the following
similar comment in the past:
{precedent}
The moderator made a decision of "{decision}".
Answer with "relevant" if the past comment should
be considered in the decision for the new comment, or
"not relevant" if it should not.
A.3. Prompts for Constitutional Grounding
Prompt for rule determination:
Below are a list of guidelines for making your decision:
{instructions}
The content you are deciding is:
{input}
Assess whether each rule applies, and then make your
final decision.
Think step by step.
Appendix B Screening Survey
B.1. Screening Questions for toxicity
These questions are adapted from a subset of questions taken from survey questions annotators had to answer in the original toxicity perspectives dataset (Kumar et al., 2021).
  1. (1)
    What types of sites do you use regularly?
    • Social Networking (Facebook, Twitter)
    • Video (YouTube, Twitch)
    • News (CNN, Fox, NYT, WSJ)
    • Community Forums (Reddit, Craigslist, 4chan)
    • Email or messaging (Gmail, WhatsApp, Facebook Chat)
  2. (2)
    For the sites you use, have you ever seen posts or comments you consider toxic?
    • Yes
    • No
  3. (3)
    Have you ever personally been the target of toxic posts or comments?
    • Yes
    • No
Questions 2 and 3 were used to determine which group the annotator was placed in.
B.2. Screening Questions for mod
The subreddits used in these questions reflect the top-10 subreddits selected during the process described in subsubsection 4.1.1. Participants were only recruited for the task if they indicated that they frequented one of the 3 communities selected in 4.1.2.
  1. (1)
    Do you frequent any of the following subreddits? Check all that apply.
    • r/Coronavirus
    • r/classicwow
    • r/Games
    • r/heroesofthestorm
    • r/ShingekiNoKyoujin
    • r/AmItheAsshole
    • r/CanadaPolitics
    • r/RPClipsGTA
    • r/LabourUK
  2. (2)
    Are you a moderator in any of the aforementioned subreddits?
    • r/Coronavirus
    • r/classicwow
    • r/Games
    • r/heroesofthestorm
    • r/ShingekiNoKyoujin
    • r/AmItheAsshole
    • r/CanadaPolitics
    • r/RPClipsGTA
    • r/LabourUK